INFRARED PHOTODISSOCIATION SPECTROSCOPY OF PROTONATED WATER CLUSTERS NEAR THE MAGIC NUMBER H^{+}(H_{2}O)_{21}

\underline{R. S. WALTERS}, T. D. JAEGER, M. A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA, 30602-2556; J. W. SHIN, N. I. HAMMER, E. G. DIKEN, M. A. JOHNSON, Sterling Chemistry Laboratory, Yale University, New Haven, CT, 06520; R. A. CHRISTIE and K. D. JOR-DAN, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA, 15260.

H^{+}(H_{2}O)_{n} complexes in the intermediate size range (n=10-30) are studied by infrared photodissociation spectroscopy near the symmetric and asymmetric stretching vibrations of water (3657 cm^{−}^{1}, 3756 cm^{−}^{1}). The complexes fragment by the loss of intact, multiple water molecules and their infrared spectra are compared to the predictions of theory. The spectrum for the magic number n=21 cluster shows only one distinct feature in the O-H region indicating all the water molecules occupy similar binding sites. Surprisingly, a single band is also observed for n=22 and no photodissociation is observed for the calculated fundamental of the H_{3}O^{+} ion core near 2500 cm^{−}^{1}. Possible interpretations of the above results will be discussed.